OpenStudy (brookemariexx):

The surface area of a right circular cylinder of height 4 feet and radius r feet is given by S(r)=2πrh+2πr2. Find the instantaneous rate of change of the surface area with respect to the radius, r, when r = 4. 24π 16π 64π 20π

OpenStudy (brookemariexx):

@zepdrix I got some of the work done, but stuck on a step

OpenStudy (brookemariexx):

S(r) = 2pi r h + 2pi r ^2 S(r) = 2pi r * 4 + 2pi r ^2 S(r) 8 pi r + 2pi r ^2 (8 pi x)’ = 8pi(x)’ = 8 pi(1) this is what I have so far

zepdrix (zepdrix):

They said `when r=4`, it looks like you plugged in h=4 though, hmm that's no bueno :o

OpenStudy (brookemariexx):

8pi +4pi (4)^2?

zepdrix (zepdrix):

Oh I didn't read the question properly lolol :) The first line says height of 4. woops!

OpenStudy (brookemariexx):

I think it is 16? 8pi +4pi (4) = 8pi +8 pi = 16

zepdrix (zepdrix):

\[\large\rm S(r)=2\pi r h+2\pi r^2\]Plugging in 4 for h,\[\large\rm S(r)=8\pi r+2\pi r^2\]Hmm, did you forget to differentiate the second term maybe? Power rule on the r^2,\[\large\rm S'(r)=8\pi+2\cdot2\pi r\]

OpenStudy (brookemariexx):

Okay, so 20?

OpenStudy (brookemariexx):

20 pi

OpenStudy (brookemariexx):

oh wait no

OpenStudy (brookemariexx):

would it be 16pi?

zepdrix (zepdrix):

\[\large\rm S'(r)=8\pi+4\pi r\]So plug in r=4,\[\large\rm S'(4)=8\pi+4\pi (4)\]

zepdrix (zepdrix):

Ya it's probably 16.. or 20.. or something else :D

OpenStudy (brookemariexx):

haha well thank you!